From the top of a tower 40 m, a ball is projected upwards with a speed of 20 m/s at an angle of 30° with the horizontal. if g=10m/s^2, after how long will the ball hit the ground?

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
The velocity is 20m/s and the ball is thrown from a 40m tower
vertical component of velocity is vsin30 =10m/s

so 2nd law of newton 
s=ut+1/2(at2)
-40 =20t - 5t2 
Sp 5t2-20t-40=0
t2-4t-8=0
 
t= 5.5 or -1.46
so the answer is 5.5  s